7.4 An automobile manufacturer purchases primary nonferrous metals from Apex, Inc. There is an indication that Apex will need to increase capacity to meet the automobile company demands. The proposed facility will be located next to their existing plant on a site that will cost $500.000. The plant is estimated to have a $15MM capital investment. Determine the depreciation in the fifth year using the MACRS method with the half-year convention.
Answer #1

Total capital investment = 15000000+500000=15500000

According to MACRS, manufacturers of nonferrous metals falls in the category of 7-year class. The depreciation rate for the 7-year life asset for 5th year is 8.93%

Therefore,

Depreciation = 15500000*8.93%= $1384150